OpenStudy (anonymous):

2x-3y-z=-1 x-4y+3z=-11 2x+9y+3z=-13

OpenStudy (anonymous):

You can solve this using substitution as well.. it'll take a bit to work out though.. one sec..

OpenStudy (anonymous):

okayy :)

OpenStudy (anonymous):

ok so I came up with z=-173/52, y=-3/52, x=-5/4

OpenStudy (anonymous):

okay got it thank you